### v1

# Certified Branch-and-Bound MaxSAT Solving (Extended Version)

Abstract
Over the past few decades, combinatorial solvers have seen remarkable performance improvements, enabling their practical use in real-world applications. In some of these applications, ensuring the correctness of the solver’s output is critical. However, the complexity of modern solvers makes them susceptible to bugs in their source code. In the domain of satisfiability checking (SAT), this issue has been addressed through proof logging, where the solver generates a formal proof of the correctness of its answer. For more expressive problems like MaxSAT, the optimization variant of SAT, proof logging had not seen a comparable breakthrough until recently.
In this paper, we show how to achieve proof logging for state-of-the-art techniques in Branch-and-Bound MaxSAT solving. This includes certifying look-ahead methods used in such algorithms as well as advanced clausal encodings of pseudo-Boolean constraints based on so-called Multi-Valued Decision Diagrams (MDDs). We implement these ideas in MaxCDCL, the dominant branch-and-bound solver, and experimentally demonstrate that proof logging is feasible with limited overhead, while proof checking remains a challenge.
This is an extended version of a paper that will be published in the proceedings of AAAI 2026 Vandesande et al. (2026). It extends the published version with a technical appendix including proofs and more details.
0.1 Introduction
With increasingly efficient solvers being developed across various domains of combinatorial search and optimization, we have effectively reached a point where NP-hard problems are routinely addressed in practice. This maturation has led to the deployment of solvers in a wide range of real-world applications, including safety-critical systems and life-impacting decision-making processes—such as verifying software for transportation infrastructure Ferreira and Silva (2004), checking the correctness of plans for the space shuttle’s reaction control system Nogueira et al. (2001), and matching donors with recipients in kidney transplants Manlove and O’Malley (2014). Given these high-stakes applications, it is crucial that the results produced by such solvers are guaranteed to be correct. Unfortunately, this is not always the case: numerous reports have documented solvers producing infeasible solutions or incorrectly claiming optimality or unsatisfiability Berg et al. (2023); Brummayer et al. (2010); Cook et al. (2013); Gillard et al. (2019); Järvisalo et al. (2012).
This problem calls for a systematic solution. The most straightforward approach would be to use formal verification; that is, employing a proof assistant Bertot and Castéran (2004); Nipkow et al. (2002); Slind and Norrish (2008); de Moura and Ullrich (2021) to formally prove the correctness of a solver. However, in practice, formal verification often comes at the cost of performance Fleury (2020); Heule et al. (2022), which is precisely what has driven the success of combinatorial optimization.
Instead, we advocate for the use of certifying algorithms McConnell et al. (2011), an approach also known as proof logging in the context of combinatorial optimization. With proof logging, a solver not only produces an answer (e.g., an optimal solution to an optimization problem) but also a proof of correctness for that answer. This proof can be checked efficiently (in terms of the proof size) by an independent tool known as a proof checker. Beyond ensuring correctness, proof logging also serves as a powerful software development methodology: it supports advanced testing and debugging of solver implementations Brummayer and Biere (2009). Moreover, the generated proofs constitute an auditable trail and can be used to speed-up the generation of explanations of why a particular conclusion was reached Bleukx et al. (2026).
Proof logging was pioneered in the field of Boolean satisfiability (SAT), where numerous proof formats and proof checkers, including formally verified ones, have been developed over the years Biere (2006); Goldberg and Novikov (2003); Heule et al. (2013); Wetzler et al. (2014); Cruz-Filipe et al. (2017). For several years, proof logging has been mandatory in the main tracks of the annual SAT solving competition, reflecting the community’s strong commitment to ensuring that all SAT solvers are certifying.
In this paper, we are concerned with maximum satisfiability (MaxSAT), the optimization variant of SAT. In contrast to SAT, proof logging in MaxSAT is still relatively uncommon. While several proof systems for MaxSAT have been developed Bonet et al. (2007); Larrosa et al. (2011); Morgado and Marques-Silva (2011); Py et al. (2020) and a solver has been designed specifically to generate such proofs Py et al. (2022), only with the recent advent of the VeriPB proof system Bogaerts et al. (2023); Gocht and Nordström (2021) did a general-purpose proof logging methodology for MaxSAT see the light Vandesande et al. (2022). VeriPB, a proof system for linear inequalities over 0–1 integer variables, has since been successfully applied to MaxSAT solvers based on solution-improving search Vandesande et al. (2022); Vandesande (2023); Berg et al. (2024a) (where a SAT oracle is repeatedly queried to find solutions that improve upon the current best), as well as to solvers using core-guided search Berg et al. (2023) (where a SAT oracle is queried iteratively under increasingly relaxed optimistic assumptions). The techniques developed there have also led to certification of generalizations of MaxSAT for multi-objective problems Jabs et al. (2025).
In addition to solution-improving and ore-guided search, state-of-the-art MaxSAT solvers also employ other strategies, such as implicit hitting set and branch-and-bound search. In a companion paper Ihalainen et al. (2026), we show how to certify IHS search, and in the current paper, we develop VeriPB-based certification for branch-and-bound search Li and Manyà (2021); Li et al. (2022); Abramé and Habet (2014); Kügel (2010); Li et al. (2007); Heras et al. (2008), thereby finally covering all key search paradigm in MaxSAT solving. Modern branch-and-bound solvers combine conflict-driven clause learning (CDCL) Silva and Sakallah (1999) with a sophisticated bounding function that determines whether it is the case that no assignment refining the current node can improve upon the best solution found so far Li et al. (2022, 2021); Coll et al. (2025b). While it is well understood how to certify CDCL search using VeriPB, the main challenge lies in certifying the conclusions drawn by the bounding function. This bounding function employs look-ahead methods to generate (conditional) unsatisfiable cores: sets of literals that cannot simultaneously be true. These cores are then combined to estimate the best possible objective value that remains achievable.
In order to bring this certification to practice, there are more hurdles to overcome: state-of-the-art branch-and-bound solvers employ several other clever tricks to speed up the solving process, such as pre-processing methods as well as integrating ideas from other search paradigms. One particular technique that proved to be challenging is the use of multi-valued decision diagrams (MDDs) in order to create a CNF encoding of a solution-improving constraint (which is enabled or disabled heuristically depending on the size of the instance at hand) Bofill et al. (2020). This encoding generalizes the encoding of Abío et al. (2012) based on Binary Decision Diagrams (BDDs) by allowing splits on sets of variables instead of a single variable. While MDDs have been used in a lazy clause generation setting Gange et al. (2011), their clausal encoding has never been certified. From the perspective of proof logging, one main challenge with BDD- or MDD-based encodings is that multiple (equivalent) constraints are represented by the same variable in this encoding. In the proof, we then have to show that they are indeed equivalent. As an example, consider the constraints
$12x_{1}+5x_{2}+4x_{3}$ $\geq 6\text{\qquad and}$
$12x_{1}+5x_{2}+4x_{3}$ $\geq 9.$
Taking into account that the variables $x_{i}$ take values in $\{0,1\}$, it is not hard to see that these constraints are equivalent: all combinations of truth assignments that lead to the left-hand side taking a value at least six, must assign it a value of at least nine. In other words: the left-hand side cannot take values $6$, 77, or $8$. In general, checking whether such a linear expression can take a specific value (e.g., $7$) is well-known to be NP hard, but BDD (and MDD) construction algorithms will in several cases detect this efficiently. The main question of interest for us is how to convince a proof checker of the fact that these two constraints are equivalent, without doing a substantial amount of additional work. We achieve this using an algorithm that proves this property for all nodes in an MDD in a linear pass over its representation.
To evaluate our approach, we add proof logging to the branch-and-bound solver WMaxCDCL Coll et al. (2025b). WMaxCDCL is the current state-of-the-art in branch-and-bound MaxSAT solving, which is showcased by the fact that it won (as part of a portfolio-solver) the unweighted track of the 2024 edition and the weighted track of the 2023 edition Järvisalo et al. (2023); Berg et al. (2024b). In the experiments, we focus on evaluating the overhead in the solving time, while also measuring the time necessary to check the produced proofs; the results show that proof logging is indeed possible without significant overhead for most cases, while proof checking overhead can be improved.
The rest of this paper is structured as follows. In Section 0.2, we recall some preliminaries about MaxSAT solving and VeriPB-based proof logging. Section 0.3 is devoted to presenting the core of the MaxCDCL algorithm, focusing on look-ahead–based bounding, as well as explaining how to get a certifying version of this. In Section 0.4, we explain how BDDs and MDDs are used to encode PB constraints and how this can be made certifying. Section 0.5 contains our experimental analysis and Section 0.6 concludes the paper.
Formal details and proofs are can be found in the appendices of this document. First, the appendix it contains a short discussion on how to use the redundance-based strengthening rule for proving reification and proofs by contradiction. Next, for Section 0.3 on Certifying Branch-and-Bound with Look-Ahead, it includes the proof for Theorem 5, together with a more detailed example that shows the look-ahead call that leads to the discovery of the local cores for which clause $C_{\mathcal{Q}}$ are derived. For Section 0.4 on Certifying CNF Encodings based on Binary (and Multi-Valued) Decision Diagrams, it contains more information on the MDD encoding, together with all the formal details on how to certify MDD-based encodings for the solution-improving constraint. We also include the proofs that were omitted in the paper due to space restrictions. Importantly, the proofs presented in this technical appendix are of constructive nature on purpose: it shows exactly how to perform the derivation in the VeriPB proof system, showing how it is implemented in the solver.
0.2 Preliminaries
We first recall some concepts from pseudo-Boolean optimization and MaxSAT solving. Afterwards, we introduce the VeriPB proof system. A more complete exposition can be found in previous work Li and Manyà (2021); Bogaerts et al. (2023).
0.2.1 Pseudo-Boolean Optimization and MaxSAT
[…]
A (partial) assignment $\alpha$ is a (partial) function from the set of variables to $\{0,1\}$; it is extended to literals in the obvious way, and we sometimes identify an assignment with the set of literals it maps to $1$. We write $C\!\!\upharpoonright\!\!_{\alpha}$ for the constraint obtained from $C$ by substituting all assigned variables $x$ by their assigned value $\alpha(x)$. A constraint $C$ is satisfied under $\alpha$ if $\sum_{\alpha(\ell_{i})=1}w_{i}\geq A$. A formula $F$ is satisfied under $\alpha$ if all of its constraints are. We say that $F$ implies $C$ (and write $F\models C$) if all assignments that satisfy $F$ also satisfy $C$. A constraint $C$ (unit) propagates a literal $x$ to a value in $\{0,1\}$ under a (partial) assignment $\alpha$ if assigning $x$ the opposite value makes the constraint $C\!\!\upharpoonright\!\!_{\alpha}$ unsatisfiable.
A pseudo-Boolean optimization instance consists of a formula $F$ and a linear term $\mathcal{O}=\sum_{i}v_{i}b_{i}$ (called the objective) to be minimized, where $v_{i}$ are integers (without loss of generality assumed to be positive here) and $b_{i}$ are different literals, which we will refer to as objective literals. If $\ell$ is a literal, we write $w_{\mathcal{O}}(\ell)$ for the weight of $\ell$ in $\mathcal{O}$, i.e., $w_{\mathcal{O}}(b_{i})=v_{i}$ for all $i$ and $w_{\mathcal{O}}(\ell)=0$ for all other literals. In line with previous work on certifying MaxSAT solvers Berg et al. (2023); Vandesande (2023); Berg et al. (2024a), in this paper, we view MaxSAT as the special case of pseudo-Boolean optimization where $F$ is a conjunction of clauses. A discussion on the equivalence of this view and the more clause-centric view using hard and soft clauses is given by for example Leivo et al. (2020).
0.2.2 The VeriPB Proof System
For a pseudo-Boolean optimization instance $(F,\mathcal{O})$, the VeriPB proof system Bogaerts et al. (2023); Gocht and Nordström (2021) maintains a proof configuration $\mathcal{F}$, which is a set of constraints derived so far, and is initialized with $F$. We are allowed to update the configuration using the cutting planes proof system Cook et al. (1987):
Literal Axioms:
For any literal, we can add $\ell_{i}\geq 0$ to $\mathcal{F}$.
Linear Combination:
Given two PB constraints $C_{1}$ and $C_{2}$ in $\mathcal{F}$, we can add a positive linear combination of $C_{1}$ and $C_{2}$ to $\mathcal{F}$.
Division:
Given the normalized PB constraint $\sum_{i}w_{i}\ell_{i}\geq A$ in $\mathcal{F}$ and a positive integer $c$, we can add the constraint $\sum_{i}\lceil w_{i}/c\rceil\ell_{i}\geq\lceil A/c\rceil$ to $\mathcal{F}$.
Saturation:
Given the normalized PB constraint $\sum_{i}a_{i}l_{i}\geq A$ in $\mathcal{F}$, we can add $\sum_{i}b_{i}l_{i}\geq A$ with $b_{i}=\min(a_{i},A)$ for each $i$, to $\mathcal{F}$.
Importantly, in many cases, it is not required to show precisely why a constraint can be derived, but the verifier can figure it out itself by means of so-called reverse unit propagation (RUP) Goldberg and Novikov (2003): if applying unit propagation until fixpoint on $\mathcal{F}\land\lnot C$ propagates to a contradiction, then we can add $C$ to $\mathcal{F}$.
VeriPB has some additional rules, which are briefly discussed below. We refer to earlier work Bogaerts et al. (2023) for more details. There is a rule for dealing with optimization statements:
Objective Improvement:
Given an assignment $\alpha$ that satisfies $\mathcal{F}$, we can add the constraint $\mathcal{O}<\mathcal{O}\!\!\upharpoonright\!\!_{\alpha}$ to $\mathcal{F}$.
The constraint added to $\mathcal{F}$ is also called the solution-improving constraint; if $0\geq 1$ can be derived from $\mathcal{F}$ after applying the objective improvement rule, then it is concluded that $\alpha$ is an optimal solution.
VeriPB allows deriving non-implied constraints with the redundance-based strengthening rule: a generalization of the RAT rule Järvisalo et al. (2012) and commonly used in proof systems for SAT. In this paper, the redundance rule is mainly used for reification: for any PB constraint $C$ and any fresh variable $v$, not used before, two applications of the redundance rule allow us to derive PB constraints that express $v\Rightarrow C$ and $v\Leftarrow C$. Moreover, the redundance rule allows us to derive implied constraints by a proof by contradiction: if we have a cutting planes derivation that shows that $\mathcal{F}\land\lnot C\models 0\geq 1$, then we can add $C$ to $\mathcal{F}$. [^1] In practice, this is mimicked by applying the redundance-based strengthening rule with an empty witness; details can be found in the supplementary material.
0.3 Certifying Branch-And-Bound with Look-Ahead
In this section, we present the working of modern branch-and-bound search for MaxSAT, specifically for the MaxCDCL algorithm. Our presentation is given with proof logging in mind (we immediately explain what is essential for each step from the proof logging perspective) and hence we often take a different perspective from the original papers introducing this algorithm Li et al. (2021); Coll et al. (2025b).
MaxCDCL combines branch-and-bound search and conflict-driven clause learning (CDCL). An overview of the algorithm is given in Algorithm 1. As can be seen, this essentially performs CDCL search (this is the branching aspect), using standard techniques such as assigning variables, unit propagation, conflict analysis, and non-chronological backtracking Silva and Sakallah (1999).
MaxCDCL maintains the objective value $v^{*}$ of the best found solution so far, which is initialized as $\infty$. However, after unit propagation the lookahead procedure is called. This procedure is where the bounding aspect happens: here sophisticated techniques are used to determine whether the current assignment can still be extended to a satisfying assignment that improves upon the current best objective value. If not, the search is interrupted and a new clause forcing the solver to backtrack is learned.
It is well known how to get VeriPB-based certification for CDCL (this is done for instance in a MaxSAT setting by Vandesande et al. (2022)), so we focus here only on how to get proof logging for the lookahead method. [^2] Obviously, our implementation with proof logging also handles the other aspects of the algorithm.
Input: CNF formula $F$, objective $\mathcal{O}$
Output: optimal assigment or UNSAT
1 $F\leftarrow$ Preprocess ($F$);
2 $\alpha\leftarrow\emptyset$; $\alpha^{*}\leftarrow\textbf{UNSAT}$;  $v^{*}\leftarrow+\infty$;
3 while true do
4 unit-propagate ($F$, $\alpha$);
5 lookahead ($F$, $\mathcal{O}$, $v^{*}$, $\alpha$);
6 if conflict detected then
7    if at root level then
8       return $\alpha^{*}$
9     Analyze conflict and backtrack;
10
11 else if all variables assigned then
12    if $\mathcal{O}\!\!\upharpoonright\!\!_{\alpha}<v^{*}$ then $\alpha^{*}\leftarrow\alpha$; $v^{*}\leftarrow\mathcal{O}\!\!\upharpoonright\!\!_{\alpha}$;
13      Backtrack
14 else
15      Decide on some unassigned variable;
16
Algorithm 1 Overview of the MaxCDCL algorithm for branch-and-bound MaxSAT solving.
Let us now focus on the lookahead procedure. Before diving into the algorithmic aspects, we formalize what it aims to compute, namely a set of weighted local cores.
Definition 1.
Let $\alpha$ be an assignment, and $(F,\mathcal{O})$ a MaxSAT instance. A weighted local core of $(F,\mathcal{O})$ relative to $\alpha$ is a triple
$q=\langle w,R,K\rangle$
such that $R\subseteq\alpha$ and $K$ contains only negations of objective literals and
$F\land R\land K\models\bot,$
where $\bot$ denotes the trivially false constraint $0\geq 1$.
In other words, a local core guarantees that, given the assignments in $R$ (which are also called the reasons of $K$), the objective literals in $K$ cannot all be false, and hence at least one of the objective literals in $K$ has to incur cost.
Now, we are not just interested in finding a single local core, but in a compatible set of such cores.
Definition 2.
Let $\mathcal{Q}$ be a set of weighted local cores of $(F,\mathcal{O})$ relative to $\alpha$. We call $\mathcal{Q}$ $\mathcal{O}$-compatible, if for each objective literal $\ell$,
${\mathop{\sum\mathrlap{w}}_{\langle w,R,K\rangle\in\mathcal{Q}\land\overline{\ell}\in K}}\leq w_{\mathcal{O}}(\ell).$
In other words, $\mathcal{Q}$ is $\mathcal{O}$-compatible if the total weight of all cores containing a literal does not exceed the weight of the underlying objective literal in the objective. In what follows, if $\ell$ is an objective literal, we write $\mathit{res}(\ell,\mathcal{Q})$ for
$w_{\mathcal{O}}(\ell)-{\mathop{\sum\mathrlap{w}}_{\langle w,R,K\rangle\in\mathcal{Q}\land\overline{\ell}\in K}}$
and call this the residual weight of $\ell$ with respect to $\mathcal{Q}$. The total weight of a core set, written $w_{\mathcal{O}}(\mathcal{Q})$ is $\sum_{\langle w,R,K\rangle\in\mathcal{Q}}w$. The following proposition now formalizes how a set of weighted local cores can be used during search.
Proposition 3.
Let $\mathcal{Q}$ be an $\mathcal{O}$-compatible set of weighted local cores and assume $\beta$ is any total assignments that refines $\alpha$ and satisfies $F$.
  1. 1.
    If $w_{\mathcal{O}}(\mathcal{Q})\geq v^{*}$, then $\mathcal{O}\!\!\upharpoonright\!\!_{\beta}\geq v^{*}$.
  2. 2.
    If for some unassigned objective literal $\ell$,
    $\mathit{res}{(\ell,\mathcal{Q})}+w_{\mathcal{O}}(\mathcal{Q})\geq v^{*},$
    then it holds that if $\mathcal{O}\!\!\upharpoonright\!\!_{\beta}<v^{*}$, then $\beta(\ell)=0$.
The first case of this proposition is known as a soft conflict Coll et al. (2025a). This means that if the total weight of the core set exceeds $v^{*}$, then we are sure that all assignments that refine the current one will have an objective value that does not improve upon the best found so far. In this case, we know that the current node of the search tree is hopeless and we can backtrack (in fact, MaxCDCL will learn a clause explaining this soft conflict, as discussed later). The second case in Proposition 3 was called hardening Coll et al. (2025a). The setting captured here is that if we were to set a literal $\ell$ to be cost-incurring, then the condition in the first item would trigger. As such, we can safely propagate $\ell$ to be non-cost-incurring.
Pseudo-code of the lookahead procedure, and details as to how it finds those local cores can be found in Algorithm 2. It initializes a local assignment $\lambda$ to be equal to $\alpha$, and maintains a partially constructed core $K$, initialized as the empty set. In line 2 it initializes $\mathcal{Q}$ to be the set of all trivial weighted local cores (by going over all objective literals that are already cost-incurring in $\alpha$). In the main loop at line 2, it iteratively assigns an objective literal to be non cost-incurring and runs unit propagation. If this results in another objective literal to be propagated to be cost-incurring (line 2) or if a conflict is found (line 2) , this means we have found a local core. In principle taking $R$ equal to $\alpha$ and $K$ as constructed so far could be used for a local core. Instead, however, line 2 uses standard conflict analysis methods to further reduce this. The local core that is found in this way is added to $\mathcal{Q}$ (line 2) and $K$ and $\lambda$ are reset (line 2).
If one of the two conditions of Proposition 3 is satisfied, a new clause is learned (based on the set of local cores and standard conflict analysis techniques) and added to $F$ (lines 22). Here, we use the notation
$C_{\mathcal{Q}}:=\bigvee_{\langle w,R,K\rangle\in\mathcal{Q}\land\ell\in R}\overline{\ell}.$
Input: Formula $F$, objective $\mathcal{O}$, upper bound $v^{*}$, assignment $\alpha$
1 $\lambda\leftarrow\alpha$; $\mathcal{Q}\leftarrow\emptyset$;
2 $K\leftarrow\emptyset$;
3 for each objective literal $\ell$ s.t. $\alpha(\ell)=1$ do
4 $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{\langle w_{\mathcal{O}}(\ell),\{\ell\},\{\overline{\ell}\}\rangle\}$
5while some $\ell$ with $\mathit{res}(\ell,\mathcal{Q})>0$ is unassigned do
6 $\lambda(\ell)\leftarrow 0$; $K\leftarrow K\cup\{\overline{\ell}\}$;
7 unit-propagate ($F$, $\lambda$);
8 if some $\ell^{\prime}$ with $\mathit{res}(\ell^{\prime},\mathcal{Q})>0$ is propagated then
9    $K\leftarrow K\cup\{\overline{\ell}^{\prime}\}$;
10 else if no conflict was derived then
11    continue
12 $R,K\leftarrow\texttt{improve-core}(K)$;
13 $w\leftarrow\min_{\overline{\ell}\in K}\mathit{res}(\ell,\mathcal{Q})$;
14 $\mathcal{Q}\leftarrow\mathcal{Q}\cup\{\langle w,R,K\rangle\}$;
15 if $w_{\mathcal{O}}(\mathcal{Q})\geq v^{*}$ then break;
16 $K\leftarrow\emptyset$; $\lambda\leftarrow\alpha$
17if $w_{\mathcal{O}}(\mathcal{Q})\geq v^{*}$ then
18 $F\leftarrow F\cup\{\texttt{analyze}(C_{\mathcal{Q}})\}$;
19
20else
21 for each $\ell$ with $\mathit{res}(\ell,\mathcal{Q})+w_{\mathcal{O}}(\mathcal{Q})\geq v^{*}$ do
22    $F\leftarrow F\cup\{\texttt{analyze}(C_{\mathcal{Q}}\vee\overline{\ell})\}$ ;
23    unit-propagate ($F$, $\alpha$);
24
Algorithm 2 Overview of the lookahead procedure.
We now turn our attention to proof logging for this lookahead procedure. It consists of two important components. On the one hand, we need proofs for the core discovery (lines 22) and on the other hand, the proof that the clause added to $F$ can indeed be derived. We now show that this can indeed be done efficiently.
Proposition 4.
If $\langle w,R,K\rangle$ is a local core found by Algorithm 2, then the clause
$C_{q}:=\bigvee_{\ell\in R\cup K}\overline{\ell}.$
is RUP wrt $F$.
Proof sketch.
This follows from the fact that Algorithm 2 finds a local core $q$ after the propagation of the negation of $C_{q}$ leads to a contradiction. ∎
Theorem 5.
Let $\mathcal{Q}$ be an $\mathcal{O}$-compatible set of weighted local cores and let $|\mathcal{O}|$ be the number of objective literals.
  1. 1.
    If $w_{\mathcal{O}}(\mathcal{Q})\geq v^{*}$, then there is a cutting planes derivation that derives $C_{\mathcal{Q}}$ from the constraint $\mathcal{O}\leq v^{*}-1$ and the constraints $C_{q}$ for every $q\in\mathcal{Q}$ using at most $3|\mathcal{O}|+2|\mathcal{Q}|+1$ steps.
  2. 2.
    If for some unassigned objective literal $\ell$, $\mathit{res}{(\ell,\mathcal{Q})}+w_{\mathcal{O}}(\mathcal{Q})\geq v^{*},$ then there is a cutting planes derivation that derives $C_{\mathcal{Q}}\lor\overline{\ell}$ from the constraint $\mathcal{O}\leq v^{*}-1$ and the constraints $C_{q}$ for every $q\in\mathcal{Q}$ using at most $3|\mathcal{O}|+2|\mathcal{Q}|-2$ steps.
Note that since the clauses that are finally learned by MaxCDCL in lines 2 and 2 of Algorithm 2 are derived by standard CDCL on the clauses derived in Theorem 5, they can be proven by RUP. Moreover, when line 1 of Algorithm 1 finds a conflict at root level, the clause $0\geq 1$ is also provable by RUP, which proves the optimality of the last found solution due to the objective improvement rule.
We end this section with an example illustrating the cutting planes derivation of the first case of Theorem 5.333A more advanced example, as well as the proof for Theorem 5 can be found in the appendix.
Example 6.
Let us consider a MaxSAT instance expressed as a PBO instance $(F,\mathcal{O})$, with $\mathcal{O}=3x_{1}+5x_{2}+5x_{3}+6x_{4}$. Assume that the best found solution so far has objective value $v^{*}=7$, that the current partial assignment assigns $y_{1}$ and $y_{2}$ to true, and that algorithm 2 has found a set of local cores $\mathcal{Q}=\{q_{1},q_{2}\}$, with
$q_{1}=\langle 3,\{y_{1}\},\{\overline{x_{1}},\overline{x_{2}}\}\rangle\text{,\qquad and}$ $q_{2}=\langle 5,\{y_{2}\},\{\overline{x_{3}},\overline{x_{4}}\}\rangle\text{.}$
We can observe that this is indeed an $\mathcal{O}$-compatible set and that the first case of Theorem 5 is applicable. Hence, we will derive the clause $C_{\mathcal{Q}}$, which is $\overline{y_{1}}+\overline{y_{2}}\geq 1$.
Following Proposition 4, we can derive the clauses
$C_{q_{1}}$ $:=\overline{y_{1}}+x_{1}+x_{2}\geq 1\text{, and}$
$C_{q_{2}}$ $:=\overline{y_{2}}+x_{3}+x_{4}\geq 1$
by Reverse Unit Propagation. Multiplying $C_{q_{1}}$ and $C_{q_{2}}$ by their respective weight and adding them up results in
$3\overline{y_{1}}+3x_{1}+3x_{2}+5\overline{y_{2}}+5x_{3}+5x_{4}\geq 8$ (1)
Now, from the solution improving constraint, we can obtain
$3x_{1}+3x_{2}+5x_{3}+5x_{4}\leq 6$ (2)
by the addition of literal axioms $x_{4}\geq 0$ and $x_{2}\geq 0$ multiplied by $2$. Addition of (1) and (2), and simplification gives
$3\overline{y_{1}}+5\overline{y_{2}}\geq 2$
Dividing this by $5$ now yields $C_{\mathcal{Q}}$.
0.4 Certifying CNF Encodings based on (Multi-Valued) Decision Diagrams
When a solution is found (at line 1 of Algorithm 1), MaxCDCL decides heuristically whether or not to add a CNF-encoding of the solution-improving constraint $\mathcal{O}\leq v^{*}-1$. For many CNF encodings of pseudo-Boolean constraints, it is well-known how to get certification Vandesande (2023); Gocht et al. (2022); Berg et al. (2024a). One notable exception is the CNF encoding based on binary decision diagrams (BDD), which Bofill et al. (2020) recently generalized to so-called multi-valued decision diagrams (MDD) and is used by MaxCDCL. Most of the interesting questions from a proof logging perspective already show up for BDDs, so to keep the presentation more accessible, we focus on that case and afterwards briefly discuss how this generalizes to MDDs. Formal details and proofs are included in the supplementary material.
A Binary Decision Diagram (BDD) is a (node- and edge-)labeled graph with two leaves, labeled true ($\mathbf{t}$) and false ($\mathbf{f}$), respectively, where each internal node is labeled with a variable and has two outgoing edges, labeled true ($\mathbf{t}$) and false ($\mathbf{f}$), respectively. We write $\mathrm{child(\eta,\mathbf{f})}$ and $\mathrm{child(\eta,\mathbf{t})}$ for the children following the edge with labels $\mathbf{f}$ and $\mathbf{t}$, respectively. Each node $\eta$ in a BDD represents a Boolean function: the true and false leaf nodes represent a tautology and contradiction respectively; if $\eta$ is a node labeled $x$ with true child $\eta_{\mathbf{t}}$ and false child $\eta_{\mathbf{f}}$, it maps any (total) assignment $\alpha$ to $\eta(\alpha)$, which is defined as $\eta_{\mathbf{t}}(\alpha)$ if $x\in\alpha$ and as $\eta_{\mathbf{f}}(\alpha)$ if $\overline{x}\in\alpha$.
A BDD is ordered if there is a total order of the variables such that each path through the BDD respects this order and it is reduced if two conditions hold: (1) no node has two identical children and (2) no two nodes have the same label, $\mathbf{t}$-child and $\mathbf{f}$-child. For a fixed variable ordering, each Boolean function has a unique ordered and reduced representation as a BDD, i.e., ordered and reduced BDDs form a canonical representation of Boolean functions.
When using BDDs to encode a solution-improving constraint $\mathcal{O}\leq v^{*}-1$ as clauses, first a BDD representing this constraint is constructed, then a set of clauses is generated from this BDD.
Phase 1: Construction of a BDD
The standard way to create a reduced and ordered BDD for $\sum_{i=1}^{n}v_{i}b_{i}\leq v^{*}-1$ is to first recursively descend and create BDDs $\eta_{\mathbf{t}}$ for $\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}-a_{1}$ and $\eta_{\mathbf{f}}$ for $\sum_{i>1}^{n}v_{i}b_{i}\leq v^{*}$ and then combine them into the node $\mathrm{bdd}(x_{1},\eta_{\mathbf{t}},\eta_{\mathbf{f}})$. Using memoization, the BDD is always kept reduced: if a node is being created with the same label, and the same two children as an existing node, the existing node is returned instead; if $\eta_{\mathbf{t}}=\eta_{\mathbf{f}}$, then $\eta_{\mathbf{t}}$ is returned.
This procedure, which first creates the two children, will however always take exponential time (in terms of the number of objective literals), while this can in many cases be avoided. Crucially, all PB constraints for which a BDD node are created are of the form $\sum_{i\geq k}^{n}v_{i}b_{i}\leq A$ and for a fixed $k$, the fact that we work with Boolean variables means that many different right-hand-sides will result in an equivalent constraint. Moreover, such values can be computed while constructing the BDD using a dynamic programming approach. In the following proposition, which formalizes the construction, we say that $[l,u]$ is a degree interval for $\sum_{i\geq k}^{n}v_{i}b_{i}$ when we mean that for all values $A$ in the (possibly unbounded) interval $[l,u]$, $\sum_{i\geq k}^{n}v_{i}b_{i}\leq A$ is equivalent to $\sum_{i\geq k}^{n}v_{i}b_{i}\leq u$.
Proposition 7 (Abío et al. (2012)).
If $l_{\mathbf{t}},u_{\mathbf{t}}]$ and $\l _{\mathbf{f}},u_{\mathbf{f}}]$ are degree intervals for $\sum_{i=k+1}^{n}v_{i}b_{i}$, then
$\max(l_{\mathbf{t}}+v_{k},l_{\mathbf{f}}),\min(u_{\mathbf{t}}+v_{k},u_{\mathbf{f}})]$
is a degree interval for $\sum_{i=k}^{n}v_{i}b_{i}$.
The dynamic programming approach for creating ordered and reduced BDDs now consists in keeping track of this interval for each translated PB constraint and reusing already created BDDs whenever possible. From now on, we will identify a node in a BDD $\eta=\mathrm{bdd}(k,l,u)$ as the node that represents the boolean function $\sum_{i=k}^{n}v_{i}b_{i}\leq[l,u]$.
Phase 2: A CNF Encoding From the BDD
Given a BDD that represents a PB constraint, constructed as discussed above, we can get a CNF encoding as follows.
  • For each internal node $\eta$ in the BDD, a new variable $v_{\eta}$ is created; intuitively this variable is true only when the Boolean function represented by $\eta$ is true. In practice for the two leaf nodes no variable is created but their truth value is filled in directly. However, in the proofs below we will, to avoid case splitting pretend that a variable exists for each node.
  • For each internal node $\eta$ that is labeled with literal $b$ with children $\eta_{\mathbf{t}}=\mathrm{child(\eta,\mathbf{t})}$ and $\eta_{\mathbf{f}}=\mathrm{child(\eta,\mathbf{f})}$, the clauses
    $\overline{b}+v_{\eta_{\mathbf{t}}}+\overline{v}_{\eta}$ $\geq 1\text{, \quad and }$ (3)
    $v_{\eta_{\mathbf{f}}}+\overline{v}_{\eta}$ $\geq 1$ (4)
    are added. The first clause expresses that if $b$ is true and $\eta_{\mathbf{t}}$ is false, then so is $\eta$. The second clause expresses that whenever $\eta_{\mathbf{f}}$ is false, so is $\eta$ (this does not hold for BDDs in general, but only for the specific ones generated here, where we know $\eta$ to represent a constraint $\sum_{i\geq k}^{n}v_{i}b_{i}\leq A$ and $\eta_{\mathbf{f}}$ represents $\sum_{i\geq k+1}^{n}v_{i}b_{i}\leq A$, both with positive coefficients).
  • Finally, for the top node $\eta_{\top}$ representing $\mathcal{O}\leq v^{*}-1$, the unit clause $v_{\eta_{\top}}$ is added.
Certification
Our strategy for proof logging for this encoding follows the general pattern described by Vandesande et al. (2022), namely, we first introduce the fresh variables by explicitly stating what their meaning is in terms of pseudo-Boolean constraints and later we derive the clauses from those constraints. Due to the fact that BDDs are reduced, the first step becomes non-obvious. Indeed, the new variables represent multiple (equivalent) pseudo-Boolean constraints at once. To make this work in practice, for each node $\eta=\mathrm{bdd}(k,l,u)$, we will show we can derive the constraints that express
$v_{\eta}$ $\Rightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l,$  and $v_{\eta}\Leftarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
We will refer to these two constraints as the defining constraints for node $\eta$. These constraints together precisely determine the meaning of $v_{\eta}$, but they also contain the information that the partial sum cannot take any values between $l+1$ and $u$. Hence, proving them is expected to require substantial effort. Once we have these constraints for each node, deriving the clauses becomes very easy.
Proposition 8.
For each internal node $\eta$, the clauses (3) and (4) can be derived by a cutting planes derivation consisting of one literal axiom, one multiplication, three additions and one division from the defining constraints for $\eta$, $\mathrm{child(\eta,\mathbf{t})}$ and $\mathrm{child(\eta,\mathbf{f})}$.
Proposition 9.
Let $\eta_{\top}=\mathrm{bdd}(1,l,v^{*}-1)$ for some $l\leq v^{*}-1$ be the root node of the BDD. The unit clause $v_{\eta_{\top}}\geq 1$ can be derived by a cutting planes derivation of one addition and one deletion from the defining constraints for $\eta_{\top}$ and the solution-improving constraint.
The harder part is to actually derive the defining constraints for the internal nodes. Also this can be done in the VeriPB proof system, as we show next.
Proposition 10.
Let $\eta=\mathrm{bdd}(k,l,u)$ be an internal node with label $b_{k}$ and with children $\eta_{\mathbf{t}}=\mathrm{child(\eta,\mathbf{t})}$ and $\eta_{\mathbf{f}}=\mathrm{child(\eta,\mathbf{f})}$.
If the defining constraints for $\eta_{\mathbf{t}}$ and $\eta_{\mathbf{f}}$ are given, then the defining constraints for $\eta$ can be derived in the VeriPB proof system.
Proof Sketch.
The first step is to introduce two reification variables $v_{\eta}$ and $v_{\eta}^{\prime}$ as
$v_{\eta}\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq l\text{,\quad and}$ $v_{\eta}^{\prime}\Leftrightarrow\sum_{i\geq k}^{n}v_{i}b_{i}\leq u.$
The next step is showing that these two variables are actually the same; the difficult direction is showing that $v_{\eta}^{\prime}$ implies $v_{\eta}$. We do this by case splitting on $b_{k}$ and deriving each case by contradiction using the defining constraints for $\eta_{\mathbf{t}}$ and $\eta_{\mathbf{f}}$. To derive $\overline{x}+\overline{v}_{\eta}^{\prime}+v_{\eta}\geq 1$, assume that $b_{k}$ and $v_{\eta}^{\prime}$ are true and that $v_{\eta}$ is false. The assumption on $b_{k}$ and $v_{\eta}^{\prime}$ together with the defining constraints for $\eta_{\mathbf{t}}$ makes that $v_{\eta_{\mathbf{t}}}$ is true; the assumption that $b_{k}$ is true and $v_{\eta}$ is false together with the defining constraints for $\eta_{\mathbf{t}}$ makes that $v_{\eta_{\mathbf{t}}}$ is false. This is clearly a contradiction. Using the defining constraints of $\eta_{\mathbf{f}}$, we can in a similar way derive $b_{k}+\overline{v}_{\eta}^{\prime}+v_{\eta}\geq 1$. A short cutting planes derivation now allows us to derive that $v_{\eta}^{\prime}$ implies $v_{\eta}$ and hence derive the required defining constraints. ∎
Finally, it can happen that due to reducedness of BDDs, a node actually represents multiple constraints. The following proposition shows that also this can be proven in VeriPB.
Proposition 11.
Consider a node $\eta=\mathrm{bdd}(k,l,u)$ and let $k^{\prime}$ and $l^{\prime}$ be numbers with $k^{\prime}<k$ and $l^{\prime}=l+\sum_{i=k^{\prime}}^{k-1}v_{i}$ such that the Boolean functions $\sum_{i=k}^{n}v_{i}b_{i}\leq u$ and $\sum_{i=k^{\prime}}^{n}v_{i}b_{i}\leq u$ are equivalent.
From the defining constraints of $\mathrm{bdd}(k,l,u)$, there is a cutting planes derivation consisting of $6(k-k^{\prime})+3$ steps that derives the defining constraints for $\mathrm{bdd}(k^{\prime},l^{\prime},u)$ for reification variable $v_{\eta}$.
Generalization to MDDs
MaxCDCL not only makes use of BDDs for encoding the solution-improving constraint, but also of MDDs. The idea is as follows. In some cases, MaxCDCL can infer implicit at-most-one constraints. These are constraints of the form $\sum_{i\in I}b_{i}\leq 1$ where the $b_{i}$ are literals in the objective $\mathcal{O}$. The detection of such constraints is common in MaxSAT solvers, and certification for it has been described by Berg et al. (2023), so we will not repeat this here. Now assume that a set of disjoint at-most-one constraints has been found. In an MDD, instead of branching on single variable in each node, we will branch on a set of variables for which an at-most-one constraint has previous been derived. This means a node does not have two, but $|I|+1$ children: one for each variable in the set and one for the case where none of them is true. Otherwise, the construction and ideas remain the same. As far as certification is concerned: essentially all proofs continue to hold; the main difference is that the case splitting in the proof of Proposition 10 will now split into $|I|+1$ cases instead of two and will then use the at most one constraint to derive that the conclusion must hold in exactly one of the cases.
0.5 Experiments
We implemented proof logging in the MaxCDCL solver as described by Coll et al. (2025b). [^4] On top of the basic algorithm described above, the experiments include proof logging for a range of techniques included in MaxCDCL Coll et al. (2025b), namely clause vivification, equivalent and failed literal detection, clause subsumption and simplification due to unit propagation. Each solver call was assigned a single core on a 2x32-core AMD EPYC 9384X. The time and memory limits are enforced by runlim. [^5] https://github.com/arminbiere/runlim As benchmarks, the weighted instances of the MaxSAT Evaluation 2024 Berg et al. (2024b) are used. The implementation, together with the raw data from the experiments will be made available in the near future.
Overhead in proof logging
A scatter plot evaluating the overhead in the solving time due to proof logging is shown in Figure 1. The experiments ran with a time limit of 1 hour and a memory limit of 32GB. Out of the 353 instances that were solved without proof logging, three instances could not be solved with proof logging. Considering instances solved both with and without proof logging, the median overhead of proof logging is $23\%$, which shows that for most instances, proof logging overhead is manageable. However, we see that for $10\%$ of the instances, the overhead of proof logging is larger than $153\%$. This is in line with earlier work on certifying PB-to-CNF encodings using VeriPB, such as the work of Berg et al. (2024a). Preliminary tests indicate as possible explanation for this phenomenon that writing the defining constraints for a node in an MDD is linear in the number of objective literals. In practice, this leads to large overhead for instances with a large number of objective literals. Possible ways to circumvent this is part of planned future work.
Figure 1: Comparison of solving time with and without prooflogging.
Overhead in proof checking
While optimizing the checking time for the produced proofs was out of scope for this work, we also evaluated it by running the VeriPB proof checker VeriPB with a time limit of 10 hours and a memory limit of 64GB on the produced proofs. Out of 350 instances produced proofs, 237 were checked successfully; for the other 113 instances, the proof checker ran out of time. In the median proof verification took $23.25$ times the solving time. We believe this checking overhead to be too high for making this usable in practice. However, there are several ways in which this can be improved. On the one hand, a new proof checker is currently under development for improved performance PBOxide (2025). Simultaneously, there are ongoing efforts for building a proof trimmer for VeriPB proofs. The proofs generated by our tools can serve as benchmarks to further guide this ecosystem of VeriPB tools. On the other hand, once the proof checker is in a stable state, we can investigate which potential optimizations to the generated proofs benefit the proof checking time. Some things that come to mind are (1) using RUP with hints instead of plain RUP so that the checker does not need to perform unit propagation, and (2) deleting intermediate constraints used in the the MDD encoding to optimize memory usage of the checker. [^6] For more information on these rules, see the VeriPB repository VeriPB .
0.6 Conclusion
In this paper, we demonstrate how to add proof logging to the branch-and-bound MaxSAT solving paradigm using the VeriPB proof system. To do so, we formalize the paradigm with proof logging in mind, offering a new perspective on how these solvers operate. To put our formalization to the test, we implemented proof logging in the state-of-the-art branch-and-bound solver MaxCDCL, including the MDD encoding of the solution-improving search constraint, which is created by MaxCDCL to improve its performance. The main challenge in certifying the MDD encoding turns out to be proving that a node in an MDD represents multiple equivalent Boolean functions. Experiments show that proof logging overhead is generally manageable, while proof checking performance still leaves room for improvement.
As future work, we plan to add proof logging to literal unlocking, a recent technique by Li et al. (2025) that improves lower bounds during look-ahead.
Acknowledgments
This work is partially funded by the European Union (ERC, CertiFOX, 101122653). Views and opinions expressed are however those of the author(s) only and do not necessarily reflect those of the European Union or the European Research Council. Neither the European Union nor the granting authority can be held responsible for them.
In addition, this work is partially funded by the Fonds Wetenschappelijk Onderzoek – Vlaanderen (projects G064925N and G070521N & fellowship 11A5J25N), and by grants PID2021-122274OB-I00 and PID2024-157625OB-I00 funded by MICIU/AEI/10.13039/501100011033/FEDER, UE.